### 28 Dec 2023 v1

# On the Three Demons in Causality in Finance: Time Resolution, Nonstationarity, and Latent Factors

Abstract
Financial data is generally time series in essence and thus suffers from three fundamental issues: the mismatch in time resolution, the time-varying property of the distribution - nonstationarity, and causal factors that are important but unknown/unobserved. In this paper, we follow a causal perspective to systematically look into these three demons in finance. Specifically, we reexamine these issues in the context of causality, which gives rise to a novel and inspiring understanding of how the issues can be addressed. Following this perspective, we provide systematic solutions to these problems, which hopefully would serve as a foundation for future research in the area.
1 Introduction
Financial investments through a scientific decision making process is a major concern for both retail and professional investors. Though many literature seem to address this, yet they fail to completely explain the underlying process in a principled way. On the other hand, hedge funds and institutions give less significance to explicitly convey the reasoning behind their investment strategies. While the customers of such funds are satisfied with the profitable performances, it becomes a serious problem when they start to lose money.
To reveal logical reasons, Causality played a compelling role in uncovering highly accurate explanations of several phenomena, especially in health sciences, climate change and energy consumption. Following success in other fields, it is being actively tested in the finance domain recently. The advancements in Causal discovery (Spirtes et al., 2000) ease the procedure of arriving at approximate or better solutions. In the early days, do-calculus introduced by Pearl (2009) encouraged the researchers to transition the way of thinking with its implicit structure.
In econometrics, the factor models dominated the investment strategies for some time until the investors realized that it did not have a proper scientific foundation. Further, most of the retail traders, as well as institutions mistake that their decisions are based on the cause and effect paradigm while they are actually associational or correlational claims. Many such examples are explained through Type A and Type B Spuriosity in (Lopez de Prado, 2022). A strong theory should be able to signal any upcoming black swan events and should provide the scenarios of falsification.
Comprehending the interdependence among time series is pivotal for studying the underlying mechanisms of complex systems. The ability to delineate the causal structure within financial data holds significant potential benefits for financial agencies, particularly stakeholders in the stock market. This paper concentrates on three demons in the the causal discovery of financial data: low observed time resolution, non-stationary data, and the existence of latent factors.
Many time series causal discovery methods, such as Granger causality (Granger, 1969), typically assume that the frequency at which the data is measured aligns with the true causal frequency of the underlying process. However, numerous causal processes unfold rapidly, and the time series data we observe often has a frequency much lower than the causal frequency, usually representing a temporal aggregation of the original causal processes (Fisher, 1970). Extensive investigations have indicated that temporal aggregation can introduce errors in the estimation of causal relations (Marcellino, 1999; Breitung and Swanson, 2002; Rajaguru and Abeysinghe, 2008) This raises a fundamental question: how can we leverage low-resolution, temporally aggregated data to unveil meaningful causal relations?
Another focal point in causal discovery and finance is the issue of data nonstationarity, influencing the outcomes of causal discovery methods applied to financial data. This nonstationarity implies that causal relations and their strengths vary with time (Zhang et al., 2015). Within finance, this phenomenon is commonly referred to as concept drift, signifying the changing distribution of data over time (Lu et al., 2018). Traditional causal discovery methods such as PC (Spirtes et al., 2000), GES (Chickering, 2002), and LiNGAM (Shimizu et al., 2006) were originally designed for the i.i.d. case and did not account for distribution shifts. A recent method, PC-MCI (Runge et al., 2019), is specifically tailored for time series data; however, it still exhibits limitations when confronted with changing causal relations over time.
Moreover, in financial data analysis, latent factors refer to unobservable variables that underlie the observed market dynamics and influence multiple observable variables simultaneously (Lopez de Prado, 2022). The concept of latent factors is pivotal in causal discovery within the financial domain as it captures the latent structures and commonalities inherent in complex financial systems. The potential bad effect arises when the true causal relationships between observed variables are obscured by the influence of latent factors, introducing spurious correlations and complicating the accurate identification of causal structures (Spirtes et al., 2000). Consequently, latent factors demand careful consideration and rigorous methods in causal discovery processes to distinguish true causal relationships from those induced by the latent variables, ensuring the robustness and reliability of the discovered results in financial data analysis.
The goal of this paper is to bring attentions into the three demons in the intersection of causality and finance data analysis, presenting systematic solutions from a causal perspective. In Section 1.1, we provide the requisite background concepts in causal discovery. Section 2 delves into the issue of low resolution and temporal aggregated data, offering a meticulous mathematical analysis that demonstrates, under specific technical assumptions, the potential unveiling of causal relations through the direct application of temporal aggregated data to instantaneous causal discovery methods. In Section 3, we explain the nonstationary property of data distribution from both financial and causal perspectives, introducing our proposed method, CD-NOD, which strategically considers time as a confounding factor. Section 4 scrutinizes the adverse effects of latent causal factors, subsequently introducing our recent work involving rank deficiency methods. Our empirical findings on causal relations among select S&\&&P 100 stocks are presented in Section 5. Finally, Section 6 provides a succinct discussion and conclusion, consolidating our insights into a comprehensive framework for navigating the intricate landscape of causality and financial data analysis.
Our contribution can be summarised as below:
  • We delineate three pivotal issues concerning causality in financial data and undertake a comprehensive exploration on these challenges in a systematic manner.
  • We present well-defined methods tailored to address each of the three identified issues. Our preliminary findings reveal compelling and novel insights, providing a foundation for further exploration.
  • Our empirical study validates the effectiveness of our proposed methods in real-life scenarios in finite sample cases.
1.1 Concepts and Terminology in Causal Discovery
To furnish a thorough introduction to causal discovery methods applied to financial data in the subsequent sections, we present key concepts within the field of causal discovery below.
Structural Causal Model. The Structural Causal Model (SCM) comprises a set of endogenous variables ($V$) and a set of exogenous variables ($U$), interconnected by a set of functions ($F$) that illustrate how the changes of exogenous variables $U$ impact the values of endogenous variables $V$ in the data generation process (Pearl, 2009). For instance, for $x_{i}\in V$, $u_{i}\in U$, and $f_{i}\in F$, the model $x_{i}:=f_{i}(Pa(x_{i}),u_{i}),i=1,\dots,d$, signifies the assignment of the value $x_{i}$ to a function $f_{i}$ of its structural parents $Pa(x_{i})$ and exogenous variable $u_{i}$. For each SCM, a causal graph $\mathcal{G}$ is correspondingly generated by representing each $x_{i}$ with each vertex and adding directing edges from each parent variable in $Pa(x_{i})$ to the child $x_{i}$.
Causal Sufficiency. A set of observed variables is deemed causally sufficient if all common causes of these variables are also observed. In other words, there is no latent confounder in the causal graph and all causally related variables are observed.
Causal Markov Condition. In the causal graph, each variable is independent of its non-descendants conditional on all of its direct causes (parents). This relationship establishes a connection from the causal graph to the statistical properties.
Markov Equivalence Class. A set of causal graphs belong to the same Markov Equivalent Class if they entail the same conditional independence relations among the observed variables, though the specific causal structures may be different.
Faithfulness. A probability distribution P is faithful to a causal graph $\mathcal{G}$ if all conditional independence relations in the probability distribution P are entailed by the causal graph $\mathcal{G}$. This relationship establishes a connection from the statistical properties to the causal graph.
2 Time Resolution and Aggregation
It is widely acknowledged that causation unfolds over time, with reactions manifesting not instantly but with a slight time lag, prompting the analysis of time series data. Time series data are typically collected at fixed frequencies. However, numerous causal processes transpire too rapidly for direct observation (Fisher, 1970). The data presented to us lack the temporal granularity to capture such swift reactions; instead, what we observe is usually a temporal aggregation of the original causal processes. This situation is prevalent in the financial field. For instance, in the stock market, where changes occur at second intervals, log returns are aggregated at daily, hourly, and minute scales (Narang, 2013).
The detrimental effects of temporal aggregation on time series models have been examined in previous studies (Tiao, 1972; Weiss, 1984), and subsequent investigations have revealed that temporal aggregation can introduce errors in the estimation of causal relations (Marcellino, 1999; Breitung and Swanson, 2002; Rajaguru and Abeysinghe, 2008). Classical time series causal discovery methods, such as Granger causal analysis (Granger, 1969), exhibit sensitivity to temporal aggregation(Breitung and Swanson, 2002).
Numerous econometricians and statisticians have endeavored to recover high-frequency time series by employing disaggregation techniques on low-frequency time series. Stram and Wei (1986) conducted disaggregation by reinterpreting the generalized least squares problem as a time series task applied to a series of time series totals. Harvey and Chung (2000) established an appropriate time series model in state-space form and computed the underlying instantaneous change using the Kalman filter. Proietti (2006) advocated the application of state space methods, particularly dynamic regression models, as a means to address the challenges associated with temporal disaggregation.
While substantial efforts have been devoted to the disaggregation of time series data, the endeavor to estimate causal structures from temporally aggregated time series remains relatively limited. Fortunately, recent demonstrations indicate that, in the context of a linear causal process,when the number of raw data points forming each observed, averagely aggregated data point is sufficiently large, the observed time series exhibits behaviors akin to independence and identical distribution (i.i.d.) (instantaneous) status (Gong et al., 2017). This allows for the direct application of instantaneous causal discovery methods to shed light on underlying true time-delayed causal relations.
Let us consider such a causal process modeled by typical VAR model:
$\mathbf{X}_{t}=\mathrm{A}\mathbf{X}_{t-1}+\mathbf{e}_{t}$ (1)
where $\mathbf{X}_{t}=(X_{t,1},X_{t,2},\dots,X_{t,n})^{T}$ represents the original data vector, $\mathbf{e}_{t}=(e_{t,1},e_{t,2},\dots,e_{t,n})^{T}$ is the temporally and contemporaneously independent noise vector, and $\mathrm{A}$ is the causal transition matrix.
Furthermore, the temporally aggregated data (observed data)
$\tilde{\mathbf{X}}_{t}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{X}_{i+(t-1)k}$ (2)
is obtained by taking the average of every successive k data points. Thus, we have temporally aggregated time series $\tilde{\mathbf{X}}_{1:T}=(\tilde{\mathbf{X}}_{1},\tilde{\mathbf{X}}_{2},\dots,\tilde{\mathbf{X}}_{T})$.
Merging (1) into (2), we have:
$\tilde{\mathbf{X}}_{t}$ $=\frac{1}{k}\sum_{i=1}^{k}\mathbf{X}_{i+(t-1)k}$ (3)
$=\frac{1}{k}\sum_{i=1}^{k}\left(\mathrm{A}\mathbf{X}_{i+(t-1)k-1}+\mathbf{e}_{i+(t-1)k}\right)$
$=\mathrm{A}\left[\frac{1}{k}\sum_{i=1}^{k}\mathbf{X}_{i+(t-1)k}-\frac{1}{k}\left(\mathbf{X}_{tk}-\mathbf{X}_{(t-1)k}\right)\right]+\frac{1}{k}\sum_{i=1}^{k}\mathbf{e}_{i+(t-1)k}$
For the high-frequency causal process, it is definitely that the aggregation factor $k$ is very large. As $k\rightarrow\infty$, $\frac{1}{k}\left(\mathbf{X}_{tk}-\mathbf{X}_{(t-1)k}\right)\rightarrow 0$. Therefore, as $k\rightarrow\infty$, we have
$\tilde{\mathbf{X}}_{t}=\mathrm{A}\tilde{\mathbf{X}}_{t}+\tilde{\mathbf{e}}_{t}$ (4)
where $\tilde{\mathbf{e}}_{t}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{e}_{i+(t-1)k}$ is the mixture of independent component vectors $\mathbf{e}_{t}$, and elements in $\tilde{\mathbf{e}}_{t}$ are contemporaneous independent to each other. Clearly, $\tilde{\mathbf{X}}_{t}$ exhibits behaviors resembling a linear instantaneous causal process in the i.i.d. case, and hence instantaneous causal discovery methods can be directly applied to aggregated data to discover meaningful causal relationships.
3 Nonstationary Relationships
3.1 From the Financial Perspective
It is well-known that financial markets has a non-stationary nature. For instance, substantial evidence suggests regime shifts in the crude oil market, as highlighted by (Vo, 2009). Financial market time series data are prone to undergo unpredictable changes, which is known as concept drift (Lu et al., 2018; Žliobaitė et al., 2016; Gama et al., 2014).
Concept drift is a phenomenon where the distribution of streaming data undergoes possibly unpredictable changes over time, as described in (Lu et al., 2018). In formal terms, when the feature vector is represented by $x$ and the label by $y$, concept drift is characterized by the time-dependent alteration in the joint distribution of $p(x,y)$, as detailed in (Lu et al., 2014).
Figure 1: Scatter Plot of PFE / BA Daily Stock Prices
Figure 2: PFE / BA Daily Stock Prices
Example 1
Figures 1 and 2 demonstrate the evolving correlation between the stock price time series of Pfizer Inc. (PFE) and The Boeing Company (BA) from 2008 to 2023. This relationship has experienced significant shifts, particularly during the pre- and post-COVID periods. Between 2008 and 2019, a time of economic expansion, both PFE and BA exhibited an upward trend in stock prices. The correlation observed in this era, represented by the darkest green scatters in Figure 1, could be influenced by broader economic factors (confounders). The onset of COVID-19 in early 2020 saw a sharp decline in BA’s stock price. However, during 2020, both stocks generally increased, displaying a notably stronger correlation (evident from the medium green scatters’ slope in Figure 1). This change is possibly linked to Pfizer’s vaccine development positively impacting travel confidence, thereby affecting the aviation sector. In 2022, this correlation moderated, aligning with the evolving COVID-19 situation. By 2023, as COVID-19 concerns eased, the correlation between PFE and BA stock prices significantly weakened, approaching a negligible level (indicated by the lightest yellow scatters in Figure 1).
Conducting a linear regression on the data from different years (as shown by the varying scatter colors in Figure 1) reveals distinct slopes and intercepts for each year. The slope, though not a precise measure, suggests the causal influence of PFE on BA, while the intercept reflects BA’s price shifts attributable to other factors. These variations highlight a clear nonstationary relationship between the two companies.
To handle concept drift, prior works usually rely on the assumption that the latest data contains more useful information. Researchers have proposed to use retraining based methods (Bach and Maloof, 2008), ensemble based methods (Gomes et al., 2017), incremental adaptation based methods (Domingos and Hulten, 2000) to tackle concept drift for the general time series tasks. In recent developments, new approaches have been introduced specifically for financial applications. Li et al. (2022) suggests initially training a model to predict future data distributions. This model is then used to create training examples that mirror the anticipated future data, which are subsequently utilized to train machine learning models. On the other hand, Riva et al. (2022) advocates for treating the trading task as a non-stationary reinforcement learning problem and selects the most effective strategy from a set of reinforcement learning models.
3.2 From Causal Discovery Perspective
The profound link between financial data analysis and causal discovery was introduced in Section 1. This raises an intriguing question: how should we approach nonstationarity in the context of causal discovery? Is this a obstacle to overcome, or can it offer unexpected advantages?
On one hand, conventional causal discovery methods designed for i.i.d. samples (e.g., PC Spirtes et al. (2000) and GES (Chickering, 2002)), or other fixed functional causal-based methods (e.g., LiNGAM), will give incorrect results with nonstationarity. These methods falter because the conditional independence relationships and the functional constraints in data with shifting distributions differ from those implied by the true causal structure. Even methods tailored for time-series data (e.g., PC-MCI (Runge et al., 2019)), also fail to identify the correct causal structures due to nonstationarity affecting not just individual time series, but also the evolving causal relationships between them. A potential solution involves Bayesian change point detection to segment the data into approximately stationary time frames for causal discovery on each frame. However, this approach is limited with the assumption of non-continuous change and the focus of changing point detection is on marginal or joint distributions rather than conditional ones with causal interpretations.
On the other hand, it can be noted that concept drift in financial data (i.e., distributional change) and causal models are heavily coupled: causal models offer a compact way to describe data-generating processes (Huang et al., 2020). From a causal perspective, the concept drift often stems from alterations in a small subset of variables within their causal generating processes over time, as the joint distribution can be factorized into conditional distributions of variables’ causal generating processes, and the joint drift can be attributed to changes in the causal mechanisms of a few variables (the so-called minimal-change principle, as another way to put faithfulness). This understanding sheds light on the mechanics behind concept drift and is pivotal in pinpointing the primary factors driving these shifts. Our method, Constraint-based Causal Discovery from Heterogeneous/Nonstationary Data (CD-NOD), elaborated in Huang et al. (2020), leverages this insight to unravel these complex dynamics. More interestingly, nonstationary relationships are not necessarily obstacles to be overcome. On the contrary, they can actually provide more information to identify causal directions which might remain unidentifiable with i.i.d. data.
CD-NOD addresses the following three questions: 1. “How can we efficiently identify variables with changing local mechanisms and reliably recover the skeleton of the causal structure over observed variables?”, 2. “How can we take advantage of the information carried by distribution shifts for the purpose of identifying causal directions?”, and 3. “How can we extract from data a low-dimensional and potentially interpretable representation of changes, the so-called “driving force” of changing causal mechanisms?” (Huang et al., 2020). Regarding question 1, CD-NOD introduces a surrogate variable (which, in the case of finance data here, is the time index) to characterize hidden unknown quantities that lead to the changes over time. By involving this surrogate variable into constraint-based causal discovery methods together with all other variables, the true causal skeleton can be correctly identified, because this surrogate variable can explain all the nonstationarity, and thus by conditioning on it, the (conditional) independence relationships between observed variables are the same as those implied by the true causal structure. Regarding Question 2, the modularity property (also known as the independent change property) of causal systems (Pearl, 1988) is exploited, i.e., when there are no confounders for cause and effect, $P(\texttt{cause})$ and $P(\texttt{effect}|\texttt{cause})$ are either fixed or change independently. Regarding Question 3, note that the variables adjacent to the time index surrogate are exactly the ones that have changing causal mechanisms over time. With the causal structure and the changing causal modules available, CD-NOD finds a low-dimensional and interpretable mapping of the density of a variable given its parental variables and the surrogate index in a nonparametric way, which exactly captures the nonstationarity of the causal mechanism of this variable.
Here we provide only a general overview of CD-NOD to exploit nonstationarity for causal discovery. For more technical details, please refer to (Huang et al., 2020).
4 Modeling and Discovery of Causally Related Latent Factors
4.1 From Financial Perspective
In stock market prediction, it has been shown by Fama and French Fama and French (1996) that, their classical 3-factor model, which consists of the market factor, a factor based on market equity (small-minus-big, SMB), and a factor based on bookto-market equity (high-minus-low, HML), well summarizes the cross section of average stock returns as of the mid-1990s. It works well empirically Hou et al. (2017) though it assumes linearity relations among variables. However, it has become clearer recently that the Fama-French model fails to account for a wide array of asset pricing anomalies Hou et al. (2015).
Therefore, over the past 2 decades, researchers have been devoted to refining the 3-factor model, with prominent ones like the 4-factor model Hou et al. (2015) and the 5-factor model Fama and French (2015). This line of thought can be summarized as finding the variables that are essential for prediction but currently unknown or cannot be observed, and then incorporating them for prediction. For example, the 3-factor model Fama and French (1996) predicts the return of the $i$-th asset by $r_{t}^{i}-r_{t}^{f}=\alpha_{t}^{i}+\beta_{MRP}^{i}MRP_{t}+\beta_{SMB}^{i}SMB_{t}+\beta_{HML}^{i}HML_{t}+\epsilon_{t}^{i}$, where MRP is the market risk premium, SMB is the size premium, and HML is the value premium; based on that, the five-factor model incorporates two factors RMW, the difference between the returns on diversified portfolios with robust and weak profitability, and CMA, the difference between the returns on diversified portfolios of low and high investment firms.
Following this line of thought, it can be argued that one promising direction toward better prediction in the area would be unveiling the factors that are essential in the underlying dynamics but cannot be explained by the existing factors found by the field. We refer to these variables/factors as latent variables. It is possible that part of the latent variables relate to some existing concepts and can be measured, but we have no observation of them just because we do not know what are they. At the same time, some latent variables might relate to some new concepts in the field. We argue that, with cutting-edge causal discovery methods at hand, we are able to find the existence and cardinality of these latent variables, and even the causal relations among all the variables.
4.2 From Causal Discovery Perspective
In this paper, we delve into the fundamental task of analyzing financial data: identifying causal relationships among different stocks to determine influence patterns. This process is known as causal discovery, i.e., to recover the causal structure that delineates the causal relations among variables from observed data.
Causal discovery typically identifies structure by leveraging (conditional) independence relations. Rooted in Reichenbach’s Common Cause Principle (Hofer-Szabó et al., 1999), it posits that two variables are interdependent due to either a direct causal relation or a shared common cause, which if observed, will render them conditionally independent. Many early causal discovery methods, such as PC (Spirtes et al., 2000) and GES (Chickering, 2002), thus operate under the causal sufficiency assumption, i.e., assuming the absence of unmeasured common causes among observed variables, enabling accurate identification of the causal structure (up to an equivalence class).
However, real-world data often violate the causal sufficiency assumption, exemplified by the factor model discussed above in financial data, and also by viewing the changing quantitie that drives nonstationarity as latent variables. The presence of unmeasured confounders can lead to spurious causations when applying methods like PC and GES directly on observed data. These dependencies among variables remain unaccounted for by their confounders. To address this, approaches like Fast Causal Inference (FCI (Spirtes et al., 2000)) and its variants (Spirtes et al., 2013; Colombo et al., 2012; Akbari et al., 2021) are developed to leverage conditional independence tests to allow for causal discovery in the presence of unmeasured confounders.
Nevertheless, FCI’s limitation lies in that its results on observed variables may be too generalized and lack meaningful interpretation due to large indeterminacies (e.g., whenever it is possible for two measured variables to be confounded, it will indicate so). Moreover, it focuses solely on the causal relations among observed variables, without considering the relations among latent variables. It only allows the existence of unobserved confounders but has to assume their mutual independence. This limitation is also shared by methods using over-complete ICA-based techniques like (Hoyer et al., 2008; Salehkaleybar et al., 2020). Then, what if there are also causal relations among latent variables, where our interest also extends to? The following example provides an introduction to addressing this challenge:
Example 2
Suppose we observe four variables ${X_{1},X_{2},X_{3},X_{4}}$ as common children of a latent variable $L$. They adhere to a linear structural equation model (SEM) as follows:
[…]
$X_{1}$ $=aL+E_{1}$
$X_{2}$ $=bL+E_{2}$
$X_{3}$ $=cL+E_{3}$
$X_{4}$ $=dL+E_{4}$
where ${E_{1},E_{2},E_{3},E_{4}}$ are mutually exogenous noise components for each variable. In the absence of observable $L$, the observed variables ${X_{1},X_{2},X_{3},X_{4}}$ show no (conditional) independencies Applying PC, GES, or FCI directly will thus result in a fully connected graph, where there is no (conditional) independencies either. Then, is it possible to distinguish between the two cases by utilizing some statistical information other than (conditional) independencies? Fortunately, the non-adjacency among observed variables leaves trace as rank deficiency in their variance-covariance matrix: e.g., consider the variance-covariance matrix between $\{X_{1},X_{2}\}$ and $\{X_{3},X_{4}\}$, $\begin{bmatrix}ac&ad\\ bc&bd\end{bmatrix}\cdot\operatorname{cov}(L)$ – it is a singular matrix. In contrast, a truly fully connected graph would not exhibit such rank deficiency.
This example illustrates the so-called Tetrad constraint, exploits the assumption of linear causal relations in the structural equation model to identify latent variables. The rank-deficiency among the observed variables often indicate the existence of latent variables, which, can be understood as another way render the (conditional) independencies among observed variables, though the conditioned one is not available. Following Tetrad constraint, methods have been developed (Silva et al., 2006; Silva and Scheines, 2005; Sullivant et al., 2010; Spirtes, 2013b; Kummerfeld and Ramsey, 2016; Huang et al., 2022), albeit often under strict structural assumptions (e.g., at least two observed children for each latent variable). To mitigate these constraints, alternative methods have been developed to utilize e.g., higher-order information (non-Gaussianity) (Shimizu et al., 2009; Cai et al., 2019; Xie et al., 2020; Adams et al., 2021; Dai et al., 2022; Chen et al., 2022), matrix-decomposition (Anandkumar et al., 2013), and mixture oracles (Kivva et al., 2021). However, as a trade-off to better identification results, these methods require more parametric assumptions on the underlying causal process. In this paper, parametrically we only assume the linearity of causal relationships, and investigate what information can be acquired with only rank deficiency constraints.
4.3 Causal Discovery in the Presence of Latent Variables
Latent variables are ubiquitous and essential in machine learning. For example, pixels of an image are dependent and yet it can be argued that they are confounded by the latent underlying concept. Other examples would be causally related concepts in psychological studies: we are interested in the human hidden mental factors and yet we can only rely on questionnaires to measure the latent mental states. Generally speaking, identifying causal relations among latent variables is hard. Traditional causal discovery methods such as PC (Spirtes et al., 2000) have to assume the absence of latent confounders. The FCI (Spirtes et al., 2000) asymptotically returns the correct causal relations among measured variables, even in the presence of latent ones, but it is generally not informative: it does not try to locate causally related latent variables and cannot identify the relations that involve latent variables. Recent advances in overcomplete ICA provide a way to handle the latent causal discovery problem, but it assumes latent variables are independent, and doesn’t provide unique solutions in general.
Therefore, there is an urgent need for a computational method that can learn true causality including causally related latent variables and causal relations. Although it is well known to be an extremely hard problem, interestingly, our recent preliminary results have shown it can be reliably achieved by exploring certain distributional properties of measured variables - by properly making use of rank deficiency and GIN constraints. Specifically one can recover latent variables and involve the causal relations by making use of rank deficiency of covariance matrix (Dong et al., 2023) or testing for independence between combinations of measured variables known as GIN (Xie et al., 2020). We provide our solution as follows (more details can be found in our manuscript (Dong et al., 2023)).
Definition 1
(Latent Linear Causal Models) Suppose a directed acyclic graph where each variable $V_{i}$ is generated following a causal structural model with n+m variables as follows.
[…]
where out of n+m variables, n are observed and m are hidden, i.e., $\{{V}_{i}\}_{i=1}^{m+n}=\{{X}_{i}\}_{i=1}^{n}\cup\{{L}_{i}\}_{i=1}^{m}$, and $\varepsilon_{{V}_{i}}$ is the independent noise term.
We further have the rank faithfulness assumption for latent linear causal models as follows. A probability distribution $p$ is rank faithful to $\mathcal{G}$ if every rank constraint on a sub-covariance matrix that holds in $p$ is entailed by every linear structural model with respect to $\mathcal{G}$. It is the classical faithfulness assumption that is critical and prevalent in causal discovery (Spirtes et al., 2000; Huang et al., 2022; Dong et al., 2023); it holds generically on infinite data, as the set of values of the SCM’s free parameters for which rank is not faithful is of Lebesgue measure 0 (Spirtes, 2013a).
Goal. Given i.i.d. samples of observed variables $\mathbf{X}_{\mathcal{G}}:=\{\mathsf{X_{i}}\}_{i=1}^{n}$ generated per LLCM, we want to (i) locate all the latent variables $\mathbf{L}_{\mathcal{G}}$, determine their cardinalities, and identify the causal structure $\mathcal{G}$ over all the variables, including both latent and observed ones. To fulfill (i), we will leverage the information from the rank of the covariance matrix of observed variables.
Rank Constraint and Motivation. We propose to leverage information from the rank of the cross-covariance matrix over observed variables to unveil the underlying causal structure. The intuition lies in that t-separations are good graphical indicators and they can be inferred from rank constraints. We shall first introduce the definition of trek and t-separation, why d-separation is a special case of t-separation, as well as how t-separations are related to rank constraints. Then we will discuss why t-separations are particularly useful in the presence of latent variables.
Definition 2 (Treks (Sullivant et al., 2010))
A trek from $\mathsf{X}$ to $\mathsf{Y}$ is an ordered pair of directed paths $(P_{1},P_{2})$ where $P_{1}$ has a sink $\mathsf{X}$, $P_{2}$ has a sink $\mathsf{Y}$, and both $P_{1}$ and $P_{2}$ have the common source $\mathsf{Z}$.
Definition 3 (T-separation (Sullivant et al., 2010))
Let $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}_{\mathbf{A}}$, and $\mathbf{C}_{\mathbf{B}}$ be four subsets of $\mathbf{V}_{\mathcal{G}}$ in graph $\mathcal{G}$ (not necessarily disjoint). ($\mathbf{C}_{\mathbf{A}}$,$\mathbf{C}_{\mathbf{B}}$) t-separates $\mathbf{A}$ from $\mathbf{B}$ if for every trek ($P_{1}$,$P_{2}$) from a vertex in $\mathbf{A}$ to a vertex in $\mathbf{B}$, either $P_{1}$ contains a vertex in $\mathbf{C}_{\mathbf{A}}$ or $P_{2}$ contains a vertex in $\mathbf{C}_{\mathbf{B}}$.
Theorem 4 (Rank and T-separation (Sullivant et al., 2010))
[…]
Now we show why t-separation implied by rank is a good graphical indicator, especially in the presence of latent variables. Take the graph in Example 2 as an example. If we can observe $\mathsf{L}$ then we can directly do the CI test conditional on $\mathsf{L}$, and thus find all the structure. However, checking that CI relation is impossible, and by CI tests we can only find a clique among $\mathsf{X_{1}},\mathsf{X_{2}},\mathsf{X_{3}},\mathsf{X_{4}}$. Fortunately, t-separation implied by rank constraints can achieve a similar goal when $\mathsf{L}$ is not observed. Specifically, we can observe that $\text{rank}(\Sigma_{\{\mathsf{X_{1}},\mathsf{X_{2}}\},\{\mathsf{X_{3}},\mathsf{X_{4}}\}})=1$, which implies the existence of one variable being in the treks between $\{\mathsf{X_{1}},\mathsf{X_{2}}\}$ and $\{\mathsf{X_{3}},\mathsf{X_{4}}\}$. The rationale behind is that the t-separation of $\mathbf{A}$, $\mathbf{B}$ by $(\mathbf{C}_{\mathbf{A}},\mathbf{C}_{\mathbf{B}})$ can be deduced through rank without observing any element in $(\mathbf{C}_{\mathbf{A}},\mathbf{C}_{\mathbf{B}})$. This property makes t-separations particularly useful for unveiling the underlying causal structure when latent variables exist. Next, we formalize this intuition into the concept of atomic cover (a cover is a set of variables) as the minimal identifiable substructure in a graph, with effective cardinality of a set of covers $\mathcal{V}$ defined as $||\mathcal{V}||=|(\cup_{\mathbf{V}\in\mathcal{V}}\mathbf{V})|$.
Definition 5 (Atomic Cover)
Let $\mathbf{V}$ be a set of variables in $\mathcal{G}$ with $|\mathbf{V}|=k$, where $t$ of the $k$ variables are observed, and the remaining $k-t$ are latent. $\mathbf{V}$ is an atomic cover if $\mathbf{V}$ contains a single observed variable (i.e., $k=t=1$ ), or if the following conditions hold:
  • (i)
[…]
  • (ii)
    There exists a set of covers $\mathcal{N}$, with $||\mathcal{N}||\geq k+1-t$, such that every element in $\cup_{\mathbf{N}\in\mathcal{N}}\mathbf{N}$ is a neighbour of $\mathbf{V}$ and $(\cup_{\mathbf{N}\in\mathcal{N}}\mathbf{N})\cap(\cup_{\mathbf{C}\in\mathcal{C}}\mathbf{C})=\varnothing$.
  • (iii)
    There does not exist a partition of $\mathbf{V}=\mathbf{V_{1}}\cup\mathbf{V_{2}}$ such that both $\mathbf{V_{1}}$ and $\mathbf{V_{2}}$ are atomic covers.
The atomic cover is essential as it is the minimal identifiable substructure in a graph. Specifically, given the following Condition 6, we can strive to ensure that the rank deficiency property uniquely implies an atomic cover, and this is formally captured by Theorem 7 as follows.
Condition 6 (Basic Graphical Conditions for Identifiability)
A graph $\mathcal{G}$ satisfies the basic graphical condition for identifiability, if $\forall\mathsf{L}\in\mathbf{L}_{\mathcal{G}}$, $\mathsf{L}$ belongs to at least one atomic cover in $\mathcal{G}$ and no variable is involved in any triangle structure.
Theorem 7 (Uniqueness of Rank Deficiency)
[…]
Basically, Theorem 7 says that in certain conditions, we can build a unique connection between rank deficiency and atomic covers. Therefore, we can identify atomic covers by first ensuring the required conditions and then searching for combinations of $\mathcal{C}$ and $\mathcal{X}$ that induce rank deficiency. Next we briefly introduce how to design an algorithm that strives to fulfill these conditions, in order to cash out the rank-deficient property for identifying atomic covers in a graph, and consequently, identifying the whole causal structure over both latent and observed variables. More details can be found in our manuscript Dong et al. (2023).
Causal Discovery Algorithm. Our rank-based latent causal discovery consists of three phases.
Phase 1: Finding CI skeleton. In this phase we just follow the classical PC algorithm (Spirtes et al., 2000) to find the CI skeleton by using CI tests.
Phase 2: Finding causal clusters. In this phase, for each sub-structure that might contain latent variables, we make use of Theorem 7 to find every atomic covers in the graph. Thus we are interested in an effective search procedure to find combinations of sets of covers $\mathcal{C}$ and $\mathcal{X}$ (as defined in Theorem 7), where rank deficiency holds. Meanwhile, our procedure should fulfill all the conditions required in Theorem 7 as much as possible in order to make the found rank deficiency uniquely indicate an atomic cover in $\mathcal{G}$.
Phase 3: Refining causal clusters. In Phase 2, we strive to fulfill all required conditions such that we can correctly identify causal clusters and related structures. However, there still exist some rare cases where our search cannot ensure the requirement (i) in Theorem 7. In this situation, Phase 2 might produce a big cluster in the resulting $\mathcal{G^{\prime}}$ that should be split into smaller ones. Fortunately, the incorrect cluster will not do harm to the identification of other substructures in the graph, and thus in this phase we characterize and refine the incorrect ones.
Identifiability Theory. The overall identifiability theory of the proposed causal discovery algorithm is summarized in the following theorem.
Condition 8 (Graphical condition on colliders for identifiability)
In a latent graph $\mathcal{G}$, if there exists a set of variables $\mathbf{C}$ such that every variable in $\mathbf{C}$ is a collider of two atomic covers $\mathbf{V_{1}}$, $\mathbf{V_{2}}$, and denote by $\mathbf{A}$ the minimal set of variables that d-separates $\mathbf{V_{1}}$ from $\mathbf{V_{2}}$, then we have $|\mathbf{C}|+|\mathbf{A}|\geq|\mathbf{V_{1}}|+|\mathbf{V_{2}}|$.
Theorem 9 (Identifiability of Latent Causal Graphs)
[…]
The graph operators $\mathcal{O}_{\text{min}}$ and $\mathcal{O}_{s}$ in the above theorem are defined following (Huang et al., 2022); they are graphical operators that will not change the observational rank constraints. Roughly speaking Theorem 9 is saying that, if the required conditions are satisfied, our method can find a compact graphical representation of the underlying causal structure, with only some uncertainty about the direction of some edges.
5 Experiments
(a) Change point detection for SP100 from 2008 to 2013. The detected change point in 2008 can be explained as Global financial crisis and subprime mortgage crisis starting from 2007. The detected change point in 2009 can be explained as European debt crisis starting from 2009.
(b) Change point detection for SP100 from 2013 to 2018. The detected change point in 2013 could relate to Cypriot financial crisis in 2013. The detected change point in 2016 could relate to Chinese stock market crash in 2016.
(c) Change point detection for SP100 from 2018 to 2023. The detected change point in 2018 could relate to cryptocurrency meltdown. The change point in 2020 likely relates to Covid-19. The change point in 2022 could relate to the Russian-Ukraine war.
Figure 3: Baysian changing point detection result for SP100 from 2008 to 2023.
5.1 Baysian Change Point Detection
To utilize our rank-based latent causal discovery method we need to assume that our data is stationary; otherwise the nonstationary would induce rank deficiency constraints that cannot correctly reflect the underlying causal structure. Therefore, we first employ Baysian change point detection Fearnhead (2006); Adams and MacKay (2007); Xuan and Murphy (2007) to identify the changing points in the time series data.
Our results are shown in Figure 3, where most of the change points detected can relate to real-life financial crises or events. Specifically, the change point in 2008 can be explained as Global financial crisis and subprime mortgage crisis starting from 2007. The change point in 2009 can be explained as European debt crisis starting from 2009. The detected change point in 2016 could relate to Chinese stock market crash in 2016. The change point in 2018 could relate to cryptocurrency meltdown. The change point in 2020 likely relates to Covid-19. The change point in 2022 could relate to the Russian-Ukraine war.
The detected change points in 2018 and 2020 are significant and are well related to real-life events. Thus we split the SP100 data using these two change points and have three time series: (i) SP100 from Jan 2017 to Dec 2017, (ii) SP100 from Jan 2019 to Dec 2019, and (iii) SP100 from Jan 2021 to Dec 2021. We will refer to them as SP100-2017, SP100-2019, and SP100-2021 respectively in the following.
In the next section, we will show the causal discovery results using these three time series.
5.2 Rank-based Latent Causal Discovery
Figure 4: Discovered causal graph using a subset of SP100 stocks for 2017, 2019, and 2021, respectively. We can see that these three graphs have many things in common, and yet a big difference between 2019 and 2021 is that there are many edges between travel-related stocks and pharmaceutical stocks after Covid-19, while before Covid-19 it is not the case.
We apply our rank-based latent causal discovery method to Sp100-2017, Sp100-2019, and Sp100-2021 respectively. The result is shown in Figure 4. We can see that these three graphs have things in common. For example, there is a latent variable influencing all the pharmaceutical companies and that detected substructure is consistent among Sp-2017, Sp-2019, and Sp-2021. We also notice that for Sp-2017, we do not have latent variable for tech companies but for Sp-2019 and Sp-2021 we do.
Furthermore there is a big difference between 2019 and 2021, which can be well-explained by Covid-19. We can see that there are many edges between travel-related stocks (BKNG and BA) and pharmaceutical stocks (BMY, PFE, JNJ, MRK, GILD, ABBV, CVS, LLY, AMGN) in the result using Sp100-2021, while for Sp100-2017 and Sp100-2019 it is not the case. This is because when pharmaceutical companies announce successful vaccine trials or gain approval for a COVID-19 vaccine, investor confidence in the travel sector increases, boosting airline stocks. On the other hand, if there are setbacks in vaccine development or distribution, airline stocks may suffer due to anticipated continued travel restrictions.
6 Conclusion and Discussion
In this paper we discuss the demons in causality in finance: time resolution, nonstationarity, and latent factors. We follow a systematic causal perspective to look into these three issues and provided well-defined methods to deal with them. Our preliminary findings reveal compelling and novel insights as well as a foundation for further exploration.